# Large Learning Rates Simultaneously Achieve Robustness to Spurious Correlations and Compressibility

### v1

@@ Abstract
Robustness and resource-efficiency are two highly desirable properties for modern machine learning models. However, achieving them jointly remains a challenge. In this paper, we position high learning rates as a facilitator for simultaneously achieving robustness to spurious correlations and network compressibility. We demonstrate that large learning rates also produce desirable representation properties such as invariant feature utilization, class separation, and activation sparsity. Importantly, our findings indicate that large learning rates compare favorably to other hyperparameters and regularization methods, in consistently satisfying these properties in tandem. In addition to demonstrating the positive effect of large learning rates across diverse spurious correlation datasets, models, and optimizers, we also present strong evidence that the previously documented success of large learning rates in standard classification tasks is likely due to its effect on addressing hidden/rare spurious correlations in the training dataset.
@@ 1 Introduction
The requirement to function well in novel circumstances and being resource-efficient are two central challenges for modern machine learning (ML) systems, which are expected to perform critical functions with less resources and on smaller hardware, in the face of limited access to computational resources, environmental concerns about energy consumption, and tightening bottlenecks around computational hardware [11, 10, 22]. As ML-based technologies increasingly permeate every aspect of daily and industrial life [25], it becomes urgent to discover inductive biases that help the models address both challenges together.
Although no learner can be expected to perform well under arbitrary changes in the environment [76], they are expected to do so under “reasonable” distribution shifts – a capability commonly exhibited by numerous animal species. Such robustness to distributions that differ from the training data in principled ways have been studied under the umbrella term out-of-distribution (OOD) generalization [75]. Although scaling models and datasets, as well as targeting specific parts of this issue have shown some success [43, 7], OOD generalization remains an open problem both theoretically and in practice [13].
@@
To date, explaining the counterintuitive observation that overparameterization often improves rather than harms generalization, especially under SCs, remains to be illuminated, as does understanding when and why models become compressible [16, 3, 52, 4, 6, 58]. One key factor in this puzzle is the learning rate (LR), or step size, used during gradient-based training. Various studies have highlighted the impact of large LRs on generalization [40, 39, 50], model compressibility [4], and representation sparsity [2].
In light of these, in this paper we hypothesize that in deep neural networks, learning rate might play a pivotal role in achieving robustness to SCs without sacrificing efficiency. We then confirm this hypothesis via extensive analyses, making the following key contributions:
@@
    We establish that large LRs can simultaneously and consistently promote both compressibility and robustness specifically to SCs across a wide range of architectures, datasets, and optimizers.
@@
See Fig. 1 for an overview of our results. We now move on to a review of the relevant literature.
@@ 2 Related Work
[40] provide one of the earliest set of findings regarding the inductive bias of LRs in standard machine learning tasks, and examine how large vs. small LRs lead to qualitatively different features to be learned by the neural network. [28] point out how LRs in early training prevents the iterates from being locked into narrow valleys in the loss landscape, where the curvature in certain directions are high, to the detriment of the conditioning of the gradient covariance matrix. [39, 50] point out the importance of large LRs in early training [29]. [60] demonstrates the crucial role of spurious / opposing signals in early training, and how progressive sharpening [9, 41] of the loss landscape in the directions that pertain to the representation of these features lead to the eventual downweighting of such nonrobust features. [15] find that lottery-ticket style pruning methods present the most advantageous trade-off between performance, robustness, and compressibility. To date, no study has systematically explored how LR influences robustness to SCs or how it interacts with compressibility. See our suppl. material for an extended literature review.
@@ 3 Setup
We consider a classification setting, where the task of a model (e.g.,, a neural network) $f$ is to predict the discrete label $y\in\mathcal{Y}$, given a $d$-dimensional input $\mathbf{x}\in\mathbb{R}^{d}$. In order to assess the quality of a neural network represented by its weights $\mathbf{w}$, we consider a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto{\mathbb{R}}_{\geq 0}$, such that $\ell({y},f_{\mathbf{w}}(\mathbf{x}))$ measures the error incurred by predicting the label of $\mathbf{x}$ as $\operatorname*{arg\,max}_{j}f_{\mathbf{w}}(\mathbf{x})[j]$, when the true label is $y$. We define an unknown data distribution $\mu_{Z}$ over $\mathcal{Z}$, and a training dataset with $n$ elements, i.e.,, $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$, where each $\mathbf{z}_{i}:=(\mathbf{x}_{i},y_{i})\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{i.i.d.}}}}}{{\sim}}\mu_{Z}$. We then denote the population and empirical risks as $\mathcal{R}(\mathbf{w})\coloneqq\mathbb{E}_{\mathbf{x},y}\left[\ell(y,f_{\mathbf{w}}(\mathbf{x}))\right]$ and ^$\widehat{\mathcal{R}}(\mathbf{w})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},f_{\mathbf{w}}(\mathbf{x}_{i}))$. Unless otherwise noted, we utilize the (minibatch) stochastic gradient descent (SGD) algorithm for empirical risk minimization: $\mathbf{w}^{t+1}=\mathbf{w}^{t}-\eta\nabla_{\mathbf{w}^{t}}\frac{1}{b}\sum_{i\in\Omega^{t}}\ell(y_{i},f_{\mathbf{w}}(\mathbf{x}_{i})),$ where $\Omega^{t}$ is a randomly sampled fixed-size subset of the training set, $b:=|\Omega^{t}|$ is batch size, and $\eta$ is learning rate (LR; aka step size).
@@ 3.1 Spurious Correlations
OOD generalization describes the case whenever we have access to $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$, $\mathbf{z}_{i}:=(\mathbf{x}_{i},y_{i})\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{i.i.d.}}}}}{{\sim}}\mu^{\mathrm{train}}_{Z}$, yet we are interested in evaluating risk under a different distribution $\mathcal{R}(\mathbf{w})\coloneqq\mathbb{E}_{\mathbf{x},y\sim\mu^{\mathrm{test}}_{Z}}\left[\ell(y,f_{\mathbf{w}}(\mathbf{x}))\right]$, where $\mu^{\mathrm{train}}_{Z}\neq\mu^{\mathrm{test}}_{Z}$. This inequality in training and test distributions is called distribution shift. Some examples of distribution shift include subpopulation shift, where $\mu^{\mathrm{train}}_{Z}(y)\neq\mu^{\mathrm{test}}_{Z}(y)$, and mechanism shift, where $\mu^{\mathrm{train}}_{Z}(\mathbf{x}|y)\neq\mu^{\mathrm{test}}_{Z}(\mathbf{x}|y)$. A subtype of mechanism shift is of special importance in the literature and of particular importance for this paper: A model trained on $S$ is said to potentially be vulnerable to spurious correlations [^2] As is common in the literature, here we use the term spurious correlations to also include non-linear relationships induced by a confounder. in the training dataset when it is possible to define subsets of $\mathbf{x}$, $\mathbf{x}^{c}$ and $\mathbf{x}^{s}$, where $\mu^{\mathrm{train}}_{Z}(\mathbf{x}^{c}|y)=\mu^{\mathrm{test}}_{Z}(\mathbf{x}^{c}|y)$, and $\mu^{\mathrm{train}}_{Z}(\mathbf{x}^{s}|y)\neq\mu^{\mathrm{test}}_{Z}(\mathbf{x}^{s}|y)$. In such cases, $\mathbf{x}^{c}$ are frequently called core features (aka invariant features), and $\mathbf{x}^{s}$ are frequently called spurious features.
@@
In previous research that addressed this problem, the assumed data distributions can vary considerably [75, 53]. In this study we focus on one of the most canonical cases, previously called “perception tasks” [56] or “easy-to-learn” tasks [51], where the core features are perfectly informative with respect to the label in the training set, and spurious features imperfectly so. In such tasks, the former are construed to be more difficult to learn. We assume a separate generative model for spurious features, called bias label $b\in\mathcal{Y}$, as opposed to the class label $y\in\mathcal{Y}$ for core features, admitting the decomposition $\mu_{Z}(\mathbf{x}^{c},\mathbf{x}^{s},y,b)=\mu_{Z}(\mathbf{x}^{c}|y)\mu_{Z}(\mathbf{x}^{s}|b)\mu_{Z}(y,b)$, and while $\mu_{Z}^{\mathrm{test}}(y,b)=\mu_{Z}^{\mathrm{test}}(y)\mu_{Z}^{\mathrm{test}}(b)$, we have $\mu_{Z}^{\mathrm{train}}(y,b)\neq\mu_{Z}^{\mathrm{train}}(y)\mu_{Z}^{\mathrm{train}}(b)$. More specifically, we further assume $\mu_{Z}^{\mathrm{train}}(y=a,b=a)\gg\mu_{Z}^{\mathrm{train}}(y=a)\mu_{Z}^{\mathrm{train}}(b=a)$, where the mutual information between $y$ and $b$ in the training set presents a challenge for the learner. The value $\rho^{\mathrm{train}}:=1-\mu^{\mathrm{train}}_{Z}(y=b)$ determines the rate of bias-conflicting examples in the training dataset, which a learner can exploit to avoid utilizing spurious features. Although there does not exist a canonical definition of easy vs. difficult-to-learn features in the literature [57], they can be construed as the difficulty of estimating $\mu^{\mathrm{train}}_{Z}(y|\mathbf{x}^{c})$ vs. $\mu^{\mathrm{train}}_{Z}(b|\mathbf{x}^{s})$.
@@ 3.2 Compressibility
A frequently used metric for characterizing compressibility, especially for activations, is sparsity. Given a $d$-dimensional vector $\mathbf{z}\in{\mathbb{R}}^{d}$ it can be defined as $\mathrm{Sparsity}(\mathbf{z})=1-\|\mathbf{z}\|_{0}/d.$ Although this is an intuitive and useful metric, more nuanced notions of compressibility are desirable, e.g. when considering activation functions like sigmoid which do not produce 0 values. This is because a (deterministic or probabilistic) vector might be “summarizable” with a small subset of its elements regardless of the number of 00 entries in the vector. Therefore, we use $(q,\kappa)$-Compressibility inspired by [24]: $(q,\kappa)\mathrm{-Compressibility}(\mathbf{z})=1-\frac{\inf_{\|\mathbf{y}\|_{0}\leq\lceil\kappa d\rceil}\|\mathbf{z}-\mathbf{y}\|_{q}}{\|\mathbf{z}\|_{q}}.$ Intuitively, if a vector’s $(2,0.1)$-Compressibility is high, this means that this vector can be approximated with little error (in an $\ell_{2}$ sense) by using $0.1$ of its elements.
@@
[…]
@@ 3.3 Metrics for Representation Analysis
We use different metrics to quantify how much representations are affected by inputs belonging to different classes. The first metric we utilize is class separation $R^{2}$, where we adopt the definition of [33], who investigate this notion in relation to model generalization performance as well as representations’ transferability under different training losses: $R^{2}=1-{\bar{d}_{\mathrm{within}}}/{\bar{d}_{\mathrm{total}}}$, where $\bar{d}_{\mathrm{within}}=\sum_{k=1}^{K}\sum_{m=1}^{N_{k}}\sum_{n=1}^{N_{k}}\frac{1-\mathrm{sim}(x_{k,m},x_{k,n})}{KN_{k}^{2}}$ and $\bar{d}_{\mathrm{total}}=\sum_{k=1}^{K}\sum_{j=1}^{K}\sum_{m=1}^{N_{j}}\sum_{n=1}^{N_{k}}\frac{1-\mathrm{sim}(x_{j,m},x_{k,n})}{K^{2}N_{j}N_{k}}$, with $K$ denoting number of classes, $N_{j}$ number of samples with label $j$, and $\mathrm{sim}$ a similarity metric such as cosine similarity.
@@
While useful, class separation does not quantify sensitivity of specific neurons to classes. For this, we use class-selectivity index ($\mathrm{CSI}$) [37, 59]: $\mathrm{CSI}=\rho_{\mathrm{max}}\frac{\pi_{\mathrm{max}}-\pi_{\mathrm{-max}}}{\pi_{\mathrm{max}}+\pi_{\mathrm{-max}}+\epsilon}.$ As in prior works, we take $\pi_{\mathrm{max}}$ to be largest class-conditional activation mean for a given neuron, where the means are computed over a sample of inputs, and $\pi_{\mathrm{-max}}$ is the mean of the remaining classes. In contrast to previous use of this metric, given the prevalence of activation sparsity in our experiments, we multiply this fraction with $\rho_{\mathrm{max}}$, which corresponds to the ratio of instances belonging to the said class for which the neuron is activated. This factors in the “coverage” of a particular neuron for the instances belonging to a class, and prevents computing high class-selectivity for a neuron that fires barely at all. When computed over an unbiased test set, this metric serves to characterize a neuron’s sensitivity to core features. This is especially useful when the core features and spurious features overlap and attribution methods cannot be reliably used to investigate a specific neuron’s responsiveness to core features. We average over neurons’ class sensitivities to characterize a network’s core feature utilization. Note that in the presence of spurious feature labels the same can be computed for spurious features to produce what we call bias-selectivity index (BSI).

### v2

@@ Abstract
Robustness and resource-efficiency are two highly desirable properties for modern machine learning models. However, achieving them jointly remains a challenge. In this paper, we identify high learning rates as a facilitator for simultaneously achieving robustness to spurious correlations and network compressibility. We demonstrate that large learning rates also produce desirable representation properties such as invariant feature utilization, class separation, and activation sparsity. Our findings indicate that large learning rates compare favorably to other hyperparameters and regularization methods, in consistently satisfying these properties in tandem. In addition to demonstrating the positive effect of large learning rates across diverse spurious correlation datasets, models, and optimizers, we also present strong evidence that the previously documented success of large learning rates in standard classification tasks is related to addressing hidden/rare spurious correlations in the training dataset. Our investigation of the mechanisms underlying this phenomenon reveals the importance of confident mispredictions of bias-conflicting samples under large learning rates.
@@ 1 Introduction
The requirement to function well in novel circumstances and being resource-efficient are two central challenges for modern machine learning (ML) systems, which are expected to perform critical functions with less resources and on smaller hardware, in the face of limited access to computational resources, environmental concerns about energy consumption, and tightening bottlenecks around computational hardware [10, 9, 21]. As ML-based technologies increasingly permeate every aspect of daily and industrial life [25], it becomes urgent to determine the inductive biases that help the models address both challenges together.
Although no learner can be expected to perform well under arbitrary changes in the environment [80], they are expected to do so under “reasonable” distribution shifts – a capability commonly exhibited by numerous animal species. Such robustness to distributions that differ from the training data in principled ways have been studied under the umbrella term out-of-distribution (OOD) generalization [79]. Although scaling models and datasets, as well as targeting specific aspects of this issue have shown some success [45, 6], OOD generalization remains an open problem both theoretically and in practice [12].
@@
To date, explaining the counterintuitive observation that overparameterization often improves rather than harms generalization, especially under SCs, remains to be illuminated, as does understanding when and why models become compressible [15, 2, 55, 3, 5, 61]. One key factor in this puzzle is the learning rate (LR), or step size, used during gradient-based training. Various studies have highlighted the impact of large LRs on generalization [43, 40, 53], model compressibility [3], and representation sparsity [1]. Recent research confirms that LR remains a crucial hyperparameter in the training of multimodal foundational models [75].
In light of these, in this paper we hypothesize that in deep neural networks, learning rate plays a pivotal role in achieving robustness to SCs without sacrificing efficiency. We then confirm this hypothesis via extensive analyses, making the following key contributions:
@@
    We establish that large LRs simultaneously and consistently promote both compressibility and robustness specifically to SCs across a wide range of architectures, datasets, and training schemes.
@@
In Fig. 1 we share an overview of our results.
@@ 2 Related Work
[43] provide one of the earliest set of findings regarding the inductive bias of LRs in standard machine learning tasks, and examine how large vs. small LRs lead to qualitatively different features to be learned by the neural network. [29] point out how LRs in early training prevents the iterates from being locked into narrow valleys in the loss landscape, where the curvature in certain directions are high, to the detriment of the conditioning of the gradient covariance matrix. [40, 53] point out the importance of large LRs in early training [30]. [63] demonstrates the crucial role of spurious / opposing signals in early training, and how progressive sharpening [8, 78] of the loss landscape in the directions that pertain to the representation of these features lead to the eventual down-weighting of such non-robust features. [14] find that lottery-ticket style pruning methods present the most advantageous trade-off between performance, robustness, and compressibility. To date, no study has systematically explored how LR influences robustness to SCs or how it interacts with compressibility. See our suppl. material for an extended literature review.
@@ 3 Setup
We consider a classification setting, where the task of a model (e.g.,, a neural network) $f$ is to predict the discrete label $y\in\mathcal{Y}$, given a $d$-dimensional input $\mathbf{x}\in\mathbb{R}^{d}$. In order to assess the quality of a neural network represented by its weights $\mathbf{w}$, we consider a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto{\mathbb{R}}_{\geq 0}$, such that $\ell({y},f_{\mathbf{w}}(\mathbf{x}))$ measures the error incurred by predicting the label of $\mathbf{x}$ as $\operatorname*{arg\,max}_{j}f_{\mathbf{w}}(\mathbf{x})[j]$, when the true label is $y$. We define a data distribution $\mu_{Z}$ over $\mathcal{Z}$, and a training dataset with $n$ elements, i.e.,, $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$, where each $\mathbf{z}_{i}:=(\mathbf{x}_{i},y_{i})\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{i.i.d.}}}}}{{\sim}}\mu_{Z}$. We then denote the population and empirical risks as $\mathcal{R}(\mathbf{w})\coloneqq\mathbb{E}_{\mathbf{x},y}\left[\ell(y,f_{\mathbf{w}}(\mathbf{x}))\right]$ and $\widehat{\mathcal{R}}(\mathbf{w})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i},f_{\mathbf{w}}(\mathbf{x}_{i}))$. Unless otherwise noted, we utilize the (minibatch) stochastic gradient descent (SGD) algorithm for empirical risk minimization: $\mathbf{w}^{t+1}=\mathbf{w}^{t}-\eta\nabla_{\mathbf{w}^{t}}\frac{1}{b}\sum_{i\in\Omega^{t}}\ell(y_{i},f_{\mathbf{w}}(\mathbf{x}_{i})),$ where $\Omega^{t}$ is a randomly sampled fixed-size subset of the training set, $b:=|\Omega^{t}|$ is batch size, and $\eta$ is learning rate (LR; aka step size).
@@ 3.1 Spurious Correlations
OOD generalization describes the case whenever we have access to $S=\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$, $\mathbf{z}_{i}:=(\mathbf{x}_{i},y_{i})\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{i.i.d.}}}}}{{\sim}}\mu^{\mathrm{train}}_{Z}$, yet we are interested in evaluating risk under a different distribution $\mathcal{R}(\mathbf{w})\coloneqq\mathbb{E}_{\mathbf{x},y\sim\mu^{\mathrm{test}}_{Z}}\left[\ell(y,f_{\mathbf{w}}(\mathbf{x}))\right]$, where $\mu^{\mathrm{train}}_{Z}\neq\mu^{\mathrm{test}}_{Z}$. This difference between training and test distributions is called distribution shift. Some examples of distribution shift include subpopulation shift, where $\mu^{\mathrm{train}}_{Z}(y)\neq\mu^{\mathrm{test}}_{Z}(y)$, and mechanism shift, where $\mu^{\mathrm{train}}_{Z}(\mathbf{x}|y)\neq\mu^{\mathrm{test}}_{Z}(\mathbf{x}|y)$. A subtype of mechanism shift is of special importance in the literature and of particular importance for this paper: A model trained on $S$ is said to potentially be vulnerable to spurious correlations [^2] As is common in the literature, here we use the term spurious correlations to also include non-linear relationships induced by a confounder. in the training dataset when it is possible to define a decomposition of $\mathbf{x}=(\mathbf{x}^{c},\mathbf{x}^{s})$, such that $\mu^{\mathrm{train}}_{Z}(\mathbf{x}^{c}|y)=\mu^{\mathrm{test}}_{Z}(\mathbf{x}^{c}|y)$, and $\mu^{\mathrm{train}}_{Z}(\mathbf{x}^{s}|y)\neq\mu^{\mathrm{test}}_{Z}(\mathbf{x}^{s}|y)$. In such cases, $\mathbf{x}^{c}$ are frequently called core features (aka invariant features), and $\mathbf{x}^{s}$ are frequently called spurious features.
@@
In previous research that addressed this problem, the assumed data distributions can vary considerably [79, 56]. In this study we focus on one of the most canonical cases, previously called “perception tasks” [59] or “easy-to-learn” tasks [54], where the core features are perfectly informative with respect to the label in the training set, and spurious features imperfectly so. In such tasks, the former are construed to be more difficult to learn. We assume a separate generative model for spurious features. We define a bias label $b\in\mathcal{Y}$, as opposed to the class label $y\in\mathcal{Y}$ for core features, admitting the decomposition $\mu_{Z}(\mathbf{x}^{c},\mathbf{x}^{s},y,b)=\mu_{Z}(\mathbf{x}^{c}|y)\mu_{Z}(\mathbf{x}^{s}|b)\mu_{Z}(y,b)$, and while $\mu_{Z}^{\mathrm{test}}(y,b)=\mu_{Z}^{\mathrm{test}}(y)\mu_{Z}^{\mathrm{test}}(b)$, we have $\mu_{Z}^{\mathrm{train}}(y,b)\neq\mu_{Z}^{\mathrm{train}}(y)\mu_{Z}^{\mathrm{train}}(b)$. More specifically, we further assume $\mu_{Z}^{\mathrm{train}}(y=a,b=a)\gg\mu_{Z}^{\mathrm{train}}(y=a)\mu_{Z}^{\mathrm{train}}(b=a)$, where the mutual information between $y$ and $b$ in the training set presents a challenge for the learner. The value $\rho^{\mathrm{train}}:=1-\mu^{\mathrm{train}}_{Z}(y=b)$ determines the rate of bias-conflicting examples in the training dataset, which a learner can exploit to avoid utilizing spurious features. Although there does not exist a canonical definition of easy vs. difficult-to-learn features in the literature [60], they can be construed as the difficulty of estimating $\mu^{\mathrm{train}}_{Z}(y|\mathbf{x}^{c})$ vs. $\mu^{\mathrm{train}}_{Z}(b|\mathbf{x}^{s})$.
@@ 3.2 Compressibility
A frequently used metric for characterizing compressibility, especially for activations, is sparsity. Given a $d$-dimensional vector $\mathbf{z}\in{\mathbb{R}}^{d}$ it can be defined as $\mathrm{Sparsity}(\mathbf{z})=1-\|\mathbf{z}\|_{0}/d.$ Although this is an intuitive and useful metric, more nuanced notions of compressibility are desirable, e.g. when considering activation functions like sigma that does not output 0 values. This is because a (deterministic or probabilistic) vector might be “summarizable” with a small subset of its elements regardless of the number of entries in the vector that exactly equal 0. Therefore, we use $(q,\kappa)$-Compressibility inspired by [23]: $(q,\kappa)\mathrm{-Compressibility}(\mathbf{z})=1-\frac{\inf_{\|\mathbf{y}\|_{0}\leq\lceil\kappa d\rceil}\|\mathbf{z}-\mathbf{y}\|_{q}}{\|\mathbf{z}\|_{q}}.$ Intuitively, if a vector’s $(2,0.1)$-Compressibility is high, this means that this vector can be approximated with little error (in an $\ell_{2}$ sense) by using $0.1$ of its $d$ elements.
@@
While sparsity or $(q,\kappa)$-Compressibility have merit as compressibility metrics, given our interest in compressibility in relation to OOD generalization performance, quantifying network/parameter compressibility while taking the downstream effects of compression on performance into account is crucial. For this purpose we define $\kappa$-Prunability; given a predictor $f$ and its (unbiased) test accuracy $\mathrm{Acc}_{\mu^{\mathrm{test}}}(f)$ it is defined as $\kappa\mathrm{P}(f)={\mathrm{Acc}_{\mu^{\mathrm{test}}}(f^{(\kappa)})}\,/\,{\mathrm{Acc}_{\mu^{\mathrm{test}}}(f)}$, where $f^{(\kappa)}$ corresponds to $f$ with $\kappa\in[0,1]$ of its parameters pruned (set to 0). This measures the retention of (unbiased) test accuracy after pruning a fraction $\kappa$ of parameters. Here, we utilize structured (neuron/filter) pruning given its computationally desirable properties [41]. Instead of committing to a particular $\kappa$, we will use $\sum_{\kappa\in\mathbf{k}}\kappa\mathrm{P}(f)$, $\mathbf{k}=(0.1,0.2,\dots,0.9)$. Given our chosen metric, we will use the terms network prunability and network compressibility interchangeably throughout the rest of the text. See suppl. material for qualitatively identical results with alternative metrics.
@@ 3.3 Metrics for Representation Analysis
We use different metrics to quantify representations’ class-sensitivity. The first metric we utilize is class separation $R^{2}$, where we adopt the definition of [34], who investigate this notion in relation to model generalization performance as well as representations’ transferability under different training losses: $R^{2}=1-{\bar{d}_{\mathrm{within}}}/{\bar{d}_{\mathrm{total}}}$, where $\bar{d}_{\mathrm{within}}=\sum_{k=1}^{K}\sum_{m=1}^{N_{k}}\sum_{n=1}^{N_{k}}\frac{1-\mathrm{sim}(x_{k,m},x_{k,n})}{KN_{k}^{2}}$ and $\bar{d}_{\mathrm{total}}=\sum_{k=1}^{K}\sum_{j=1}^{K}\sum_{m=1}^{N_{j}}\sum_{n=1}^{N_{k}}\frac{1-\mathrm{sim}(x_{j,m},x_{k,n})}{K^{2}N_{j}N_{k}}$, with $K$ denoting number of classes, $N_{j}$ number of samples with label $j$, and $\mathrm{sim}$ a similarity metric such as cosine similarity.
@@
Though important for a classification task, class separation does not quantify sensitivity of specific neurons to classes. For this, we use class-selectivity index ($\mathrm{CSI}$) [38, 62]: $\mathrm{CSI}=\rho_{\mathrm{max}}\frac{\pi_{\mathrm{max}}-\pi_{\mathrm{-max}}}{\pi_{\mathrm{max}}+\pi_{\mathrm{-max}}+\epsilon}.$ As in prior works, we take $\pi_{\mathrm{max}}$ to be largest class-conditional activation mean for a given neuron, where the means are computed over a sample of inputs, and $\pi_{\mathrm{-max}}$ is the mean of the remaining classes. In contrast to previous use of this metric, given the prevalence of activation sparsity in our experiments, we scale this fraction with $\rho_{\mathrm{max}}$, which corresponds to the ratio of instances belonging to the said class for which the neuron is activated. This factors in the “coverage” of a particular neuron for the instances belonging to a class, and prevents computing high class-selectivity for a neuron that fires rarely for the members of the said class. When computed over an unbiased test set, this metric serves to characterize a neuron’s sensitivity to core features. This is especially useful when the core features and spurious features overlap and attribution methods cannot be reliably used to investigate a specific neuron’s responsiveness to core features. We average over neurons’ class sensitivities (in learned representations) to characterize a network’s core feature utilization. Note that if accessible, spurious feature labels can be used to compute the counterpart bias-selectivity index (BSI).